# Registration beyond Points: General Affine Subspace Alignment via Geodesic Distance on Grassmann Manifold

### v1

@@ 3.1 Notations
We denote scalars using italicized characters, vectors with boldface lowercase characters, matrices with boldface uppercase characters, and subspaces with blackboard bold characters. Given a $n$-dimensional vector $\mathbf{v}$, we define $\mathrm{bar}{\mathbf{v}}$ as $(n+1)$-dimensional vector by augmenting the $\mathbf{v}$ with 00 at the last element. Similarly, $\tilde{\mathbf{v}}$ is $(n+1)$-dimensional vector by augmenting $\mathbf{v}$ with $1$ and normalizing it to norm-1, which is $\tilde{\mathbf{v}}=[\frac{\mathbf{v}^{\top}}{\sqrt{1+\|\mathbf{v}\|^{2}}},\frac{1}{\sqrt{1+\|\mathbf{v}\|^{2}}}]^{\top}$. Also, $\left\|\mathbf{v}\right\|_{2}$ denotes Euclidean norm of the vector $\mathbf{v}$. Given a linear subspace $\mathbb{U}\subseteq\mathbb{R}^{n}$ with its orthonormal basis matrix $\mathbf{U}$, $\mathbf{P}_{\mathbb{U}}:=\mathbf{U}\mathbf{U}^{\top}\in\mathbb{R}^{n\times n}$ denotes the projection matrix which maps any vector $\mathbf{x}\in\mathbb{R}^{n}$ onto its projection $\mathbf{P}_{\mathbb{U}}\mathbf{x}$ within $\mathbb{U}$.
@@ 4.3 Solutions for 3D Registration Problem
where $\epsilon$ is inlier threshold and $\mathbf{1}$ refers to the indicator function which outputs the integer $1$ only when the input is a positive number. Then, we subsequently optimize the translation by minimizing the sum of the remaining term $g_{i}$, using rotation $\mathbf{R}^{*}$ obtained from the previous step:
@@
For the case of plane-to-plane alignment, the cost consists of three terms because both subspaces are 2-dimensional affine primitives. We modify the result to yield a compact form of two terms by utilizing the duality of Grassmann manifold: an element $\mathbb{W}\in\mathrm{Gr}(k,n)$ is isomorphic to $\mathbb{W}^{\perp}\in\mathrm{Gr}(n-k,n)$, which is $(n-k)$-dimensional orthogonal complements of $\mathbb{W}$. Applying the property to $\mathrm{Gr}(2,3)$, we can make two basis vectors of a plane spanned by the other plane bases using the normal vector, which is an orthogonal complement of the 3D plane:
@@ 5.3 Perspective-n-Line
Datasets: We used both synthetic and real data for the validation. For simulation, we generated the synthetic data by scattering 2D end-points within the range of $[0,640]\times[0,480]$ as in [39]. We also divide the sets into centered and uncentered cases, with the centered case having evenly distributed lines across the entire image, and the uncentered case having lines only on one quarter of the image. For both cases, we back-projected them with random depth to construct paired 3D lines. In uncentered case, 2D line segments are not uniformly distributed within the image, providing a challenging scenario prone to local minima. Increasing the outlier ratio to 80%, we randomly generated 500 sets for a single ratio, each set containing 100 pairs.
@@
Figure 6: Representative results from the synthetic PnL experiment. (a) Convergence of the bounds on the inliers for 100 correspondences with a 10% outlier ratio. (b) Runtime with respect to endpoint pixel noise and the number of correspondences.
@@ 8.3.2 Line-to-plane case
if $\text{if }\mathrm{d}_{\mathrm{Gr}}(\mathbf{R}_{m}\mathbf{n}^{i},\mathbf{d}^{i})\geq\frac{\theta}{2}\rightarrow\mathrm{d}_{\mathrm{Gr}}(\mathbf{P}_{\mathbf{R}_{0}\cdot\mathbb{B}^{i}}\mathbf{d}^{i},\mathbf{P}_{\mathbf{R}\cdot\mathbb{B}^{i}}\mathbf{d}^{i})\leq\mathrm{d}_{\mathrm{Gr}}(\mathbf{R}_{0}\mathbf{n},\mathbf{Rn})\leq\min(\frac{\pi}{2},\sqrt{3}\sigma_{r}).$ (52)
@@ 8.5 Algorithms
11:    $\mathcal{I}_{f}$, $\mathbf{R}^{*}$ = CorrRBnB($\mathcal{X}_{lin}$, $\mathcal{Y}_{lin}$, $\mathcal{I}_{i}$, $\mathcal{C}_{\mathbf{R}}$, $\epsilon_{\mathbf{R}}$)  % Algorithm 3
@@
13:    $\mathcal{I}_{f}$, $\mathbf{R}^{*}$ = FullRBnB($\mathcal{X}_{lin}$, $\mathcal{Y}_{lin}$, $\mathcal{C}_{\mathbf{R}}$, $\epsilon_{\mathbf{R}}$) % Algorithm 4
@@
15:$\mathbf{t}^{*}$ = TBnB($\mathcal{X}_{lin}$, $\mathcal{Y}_{lin}$, $\mathcal{I}_{f}$, $\mathbf{R}^{*}$, $\mathcal{C}_{\mathbf{t}}$, $\epsilon_{\mathbf{t}}$)  % Algorithm 5
@@
8:           ($\mathcal{I}_{f}$, $\mathbf{R}^{*}$) $\leftarrow$ LMOptimization($\mathcal{X}_{lin}$, $\mathcal{Y}_{lin}$, $\mathcal{I}_{f}$, $\mathbf{R}^{*}$)
@@
8:           $\mathbf{R}^{*}$leftarrow LMOptimization($\mathcal{X}_{lin}$, $\mathcal{Y}_{lin}$, $\mathbf{R}^{*}$)
@@
8:           ($e^{*}$, $\mathbf{t}^{*}$) $\leftarrow$ LMOptimization($\mathcal{X}_{lin}$, $\mathcal{Y}_{lin}$, $\mathcal{I}_{f}$, $\mathbf{R}^{*}$, $\mathbf{t}^{*}$, $e^{*}$ )
@@ 9.2 Comparison of Curves on Manifold
We now derive the length of the curve mapped from the straight line in the Euclidean parameter space to the manifold. First, we define the mapping between the two spaces as $\phi(\cdot):\mathbb{R}^{m}\rightarrow\mathrm{Gr}(k+1,n+1)$, where $m$ is the dimension of the intermediate space (e.g., $m=4$ for a 3D plane parameter $(a,b,c,d)$). Then, the mapped curve from the Euclidean embedding is derived as $\phi(t):=\phi(\mathbf{v}(t))$, where $\mathbf{v}(t)$ is the straight line on the parameter space connecting two features represented as $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$:
@@
In the case of 2D lines, the embedding is represented by the coordinate $\mathbf{v}=(a,b,c)$ from the line equation $ax+by+c=0$. Then, the corresponding projection matrix $\phi(\mathbf{v})\in\mathbb{R}^{3\times 3}$ in case of $(c<0)$ is:

### v2

@@ 3.1 Notations
[…]
@@ 4.3 Solutions for 3D Registration Problem
where $\epsilon$ is inlier threshold and $\mathbf{1}$ refers to the indicator function which outputs the integer 111 only when the input is a positive number. Then, we subsequently optimize the translation by minimizing the sum of the remaining term $g_{i}$, using rotation $\mathbf{R}^{*}$ obtained from the previous step:
@@
[…]
@@ 5.3 Perspective-n-Line
Datasets: We used both synthetic and real data for the validation. For simulation, we generated the synthetic data by scattering 2D end-points within the range of [0,640]×[0,480][0,640$\times$[0,480][ 0 , 640 ] × [ 0 , 480 ] as in [39]. We also divide the sets into centered and uncentered cases, with the centered case having evenly distributed lines across the entire image, and the uncentered case having lines only on one quarter of the image. For both cases, we back-projected them with random depth to construct paired 3D lines. In uncentered case, 2D line segments are not uniformly distributed within the image, providing a challenging scenario prone to local minima. Increasing the outlier ratio to 80%, we randomly generated 500 sets for a single ratio, each set containing 100 pairs.
@@
Figure 6: Additional results from the synthetic PnL experiment. (a) Convergence of the bounds on the inliers for 100 correspondences with a 10% outlier ratio. (b) Runtime with respect to endpoint pixel noise and the number of correspondences.
@@ 8.3.2 Line-to-plane case
$\text{if }\mathrm{d}_{\mathrm{Gr}}(\mathbf{R}_{m}\mathbf{n}^{i},\mathbf{d}^{i})\geq\frac{\theta}{2}\rightarrow\mathrm{d}_{\mathrm{Gr}}(\mathbf{P}_{\mathbf{R}_{0}\cdot\mathbb{B}^{i}}\mathbf{d}^{i},\mathbf{P}_{\mathbf{R}\cdot\mathbb{B}^{i}}\mathbf{d}^{i})\leq\mathrm{d}_{\mathrm{Gr}}(\mathbf{R}_{0}\mathbf{n},\mathbf{Rn})\leq\min(\frac{\pi}{2},\sqrt{3}\sigma_{r}).$ (52)
@@ 8.5 Algorithms
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 9.2 Comparison of Curves on Manifold
[…]
@@
[…]